Emily mixed together 9 gal. of Brand A fruit drink and 8 gal. of Brand B fruit drink which contains 48% fruit juice. Find the percent of fruit juice in Brand A if the mixture contained 30% fruit juice.
Answer by stanbon(75887)

Emily mixed together 9 gal. of Brand A fruit drink and 8 gal. of Brand B fruit drink which contains 48% fruit juice. Find the percent of fruit juice in Brand A if the mixture contained 30% fruit
-----
Equation:
fruit + fruit = fruit
(x/100)9 + 0.41*8 = 0.30(9+8)
---
Multiply thru by 100 to get:
9x + 41*8 = 30*17
9x = 182
x = 20.22% (Percent of fruit in Brand A)
===============
Cheers,
Stan H.